OpenStudy (anonymous):

The graph of a system of equations with the same slope will have no solutions. Always Sometimes Never

Directrix (directrix):

How many solutions does the following system have: y = 5x + 2 y = 5x + 6

OpenStudy (anonymous):

always

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

I do not agree with Always as the answer. Consider this system: x + y = 2 3x + 3y = 6 ----------- How many solutions does it have?

OpenStudy (anonymous):

never

Directrix (directrix):

In this case, y = 5x + 2 y = 5x + 6 there were no solutions because the lines are parallel.

Directrix (directrix):

In this case, x + y = 2 3x + 3y = 6 ----------- there were infinitely many solutions because the two equations are equations of the same line. So, they intersect at every point. And, a line has infinitely points. So, what is left for the correct answer.